KBr crystallizes in a face-centered lattice. The ionic radii of K+ and Br– ions are 133 pm and 195 pm, respectively.

Calculate the density and Packing efficiency for KBrl?
Answer #1

(fee crystal) atside of lattice 133 pm rza=q (ritra) r=195pm ſza = 2(133+195)pm a = 463.86pm. > a = 4.6386810-8 en. Density (I hope you understand this

Thanks